OpenStudy (anonymous):

Write a proportion for the following problem. Eric is making a scale drawing of the ground floor of his house. If the scale is 2 in = 5 ft, what number of inches will he use for a length of 55 ft? A. 2/5 = 55/x B. 2/55 = x/5 C. 2/5 = x/55

OpenStudy (anonymous):

@Directrix

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} in&ft \\\hline\\ 2&5\\ x&55 \end{array}\implies \cfrac{2}{x}=\cfrac{5}{55}\)

OpenStudy (jdoe0001):

notice, just line up the corresponding units vertically

OpenStudy (anonymous):

yes @jdoe0001

OpenStudy (anonymous):

would it be C ? @jdoe0001

OpenStudy (jdoe0001):

hmmm yeap, that'd work, since it yields the same equation

jimthompson5910 (jim_thompson5910):

Use jdoe0001's table to read down the columns to get \[\Large \frac{2}{x} = \frac{5}{55}\] or read across the rows to get \[\Large \frac{2}{5} = \frac{x}{55}\] The two equations are equivalent to one another. You can also flip the fractions of both sides to get other equivalent equations.